OpenStudy (anonymous):

A seagull can dive 45 meters below the surface of the water to reach a fish. It dives 11% of the total depth every 1 second. What is the depth of the seagull after 2 seconds? a. -2.475 meters b. -4.95 meters c. -9.9 meters d. -7.425 meters

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (phi):

do you know how to find 11% of 45 ?

OpenStudy (anonymous):

do you do 45/x 11/100?

OpenStudy (anonymous):

or is it the other way around

OpenStudy (phi):

just 45 times 11/100

OpenStudy (anonymous):

oh ok

OpenStudy (phi):

11% of the total depth every 1 second. that will be how far down it goes in 1 second to find how far down in 2 seconds multiply by 2 so the answer is \[ \frac{2\cdot 45\cdot 11}{100} \]

OpenStudy (phi):

it looks like they are using negative numbers to show the distance below the water surface

OpenStudy (anonymous):

ok i know the answer now. thank you